OpenStudy (anonymous):

Part A: Nadia rented a stage platform at $410 for 4 days. If she rents the same stage platform for 8 days, she has to pay a total rent of $770. Write an equation in the standard form to represent the total rent (y) that Nadia has to pay for renting the stage platform for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

@micahm

OpenStudy (micahm):

is this one question or is their more in the paragraph

OpenStudy (anonymous):

This One Question Part A B And C.

OpenStudy (micahm):

@Preetha

OpenStudy (micahm):

@undeadknight26

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

We can first find the slope, we are given two points: (4, 410) and (8, 770) Plug them in the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{770-410}{8-4}\)

OpenStudy (igreen):

Simplify that

OpenStudy (anonymous):

Okay This For Part A Right?

OpenStudy (igreen):

Yep

OpenStudy (igreen):

To write an equation we need the slope, so can you simplify that?

OpenStudy (anonymous):

I Got 360 And 4

OpenStudy (igreen):

Yes, now divide 360 / 4

OpenStudy (anonymous):

90

OpenStudy (igreen):

Yes, so the slope is 90.

OpenStudy (igreen):

Now let's plug this in with one of the points into point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y - 410 = 90(x - 4)\) Distribute 90 into the parenthesis: \(y - 410 = 90x - 360\) Add 410 to both sides, what's -360 + 410?

OpenStudy (anonymous):

50

OpenStudy (igreen):

Yep, now we have: \(y = 90x + 50\) So that's our equation

OpenStudy (anonymous):

And Thats All part a right . ? :)

OpenStudy (igreen):

Now let's change it to standard form. y = 90x + 50 Standard form is Ax + By = C So subtract 90x to both sides

OpenStudy (igreen):

No, that equation is in slope-intercept form, it's asking for the equation in Standard Form.

OpenStudy (anonymous):

Oh Okay

OpenStudy (anonymous):

-40?

OpenStudy (igreen):

No, we just subtract it to both sides.. y = 90x + 5 -90x + y = 5 So that's our equation in Standard Form

OpenStudy (anonymous):

Now Wha Part We Working On Now :)?

OpenStudy (igreen):

We just finished Part A.

OpenStudy (anonymous):

ok :)

OpenStudy (igreen):

Part B, y = 90x + 5 Function notation is just replacing 'y' with 'f(x)', so what do you get?

OpenStudy (anonymous):

f(x) = 90x+5

OpenStudy (igreen):

Yep, that takes care of Part B

OpenStudy (igreen):

For Part C, The x-axis will be labeled 'days' and the y-axis 'total rent'.

OpenStudy (anonymous):

Okay Now Part C && I Alos HAVE tROUBLES W. 3 MORE QUESTIONS BUT IF U CAN ILL GET TO THEM

OpenStudy (igreen):

It will have a slope of 90, and a y-intercept of 5.

OpenStudy (anonymous):

Thats It?

OpenStudy (igreen):

Yep

OpenStudy (igreen):

If you have more questions, close this one and open a new one.

OpenStudy (anonymous):

Okay